# Tunability of the magnetic properties in Ni intercalated transition metal dichalcogenide NbSe2

### v1

@@ Abstract
Ni0.33NbSe2 has been identified as an altermagnet with a Néel temperature of 84 K, although Ni0.21NbSe2 behaves like a nonmagnetic metal. In this work, we explore the magnetic phases of Ni-intercalated NbSe2 with stoichiometry Ni0.25NbSe2 (also represented as NiNb4Se8), which lies between the two previously studied compositions. Our results reveal a transition in the ground state from a stripe antiferromagnetic phase with Kramers degeneracy (below the critical value of Coulomb repulsion UC) to a ferromagnetic phase (above UC). In the Kramers antiferromagnetic state, the nearest-neighbour magnetic coupling is ferromagnetic between the layers and forms stripe-like patterns in the ab plane. This behavior contrasts with that of Ni0.33NbSe2 where the nearest-neighbour magnetic couplings are opposite; this can be attributed to the different Wyckoff positions of the magnetic atoms, which produce different space groups and different magnetic space groups. Spin-orbit coupling (SOC) reduces the value of UC, and the magnetic easy axis aligns along the z-axis. The strong magnetocrystalline anisotropy and the second-nearest neighbor coupling stabilize the Kramers antiferromagnetic and ferromagnetic phases with spins aligned along the z-axis over the non-collinear 120∘ phase with in-plane spin orientations. Given the triangular lattice and the closeness to both ferromagnetic and antiferromagnetic coupling, a metamagnetic transition is possible. Furthermore, a spin-glass phase may emerge if the distribution of intercalated Ni atoms is spatially inhomogeneous. Substantial modifications to the electronic structure accompany Ni intercalation in NbSe2 with respect to the pristine phase. Compared to pristine NbSe2, which features a hole pocket at the point, the stripe phase of Ni0.25NbSe2 instead exhibits an electron pocket at . This pocket at has predominantly 3z2-r2 character, and derives its main spectral weight from the Nb atoms. Ni intercalation also shifts the Van Hove singularity away from the Fermi level, thereby suppressing potential electronic instabilities.
@@ I Introduction
The tunability of magnetic and electronic phases in transition metal dichalcogenides (TMDs) doped or intercalated with ionic liquidsZhang et al. (2022), through self-intercalationWang et al. (2024), or with moleculesScheidt et al. (2015) has attracted significant research interest due to its potential for spintronic applications and the generation of new exotic phases. Intercalation of transition metals into TMDs modifies the electronic structure, leading to changes in the exchange interactions that govern the magnetic ordering. Focusing on the intercalation with magnetic transition metals, these materials exhibit diverse magnetic behaviors, including Weyl ferromagnetismSaika et al. (2022), chiral helimagnetismCao et al. (2020), altermagnetismVita et al. (2025) and spin glass statesKong et al. (2023); Maniv et al. (2021). This depends on the choice of dopant, concentration, and external factors such as strain and temperatureHe et al. (2024). The spatial distribution and local environment of the intercalate atoms plays a crucial role in defining the resultant magnetic order. The degree of hybridization between transition metal d-states and chalcogen p-states can significantly impact the strength and the nature of magnetic interactions.Liu et al. (2024) External stimuli, including electric gating, pressure, and strain engineering, offer further tunability. Strain can enhance SOC and modify superexchange interactions, leading to novel magnetic phases. Similarly, electrostatic gating has been demonstrated to control magnetic order dynamically, enabling voltage-driven phase transitions. These tunable properties make magnetic TMDs promising candidates for next-generation spintronic devices, including nonvolatile memories and logic applications.Nair et al. (2020); Shao et al. (2022)
@@
Several transition metal dichalcogenides host electronic instabilities such as charge density wave (CDW)Skolimowski et al. (2019); Wang et al. (2020); Cossu et al. (2020) and superconductivityXiao et al. (2022); Lei et al. (2020). In the ultrathin limit, NbSe2 hosts superconductivityUgeda et al. (2016); Xi et al. (2016) and CDWRivano et al. (2025), which can interplay with different layered stacking ordersCossu et al. (2024); Wang et al. (2022), self-intercalationYin et al. (2023) and can host topological phasesChiu et al. (2025). One of the relevant ingredients for the electronic instabilities in NbSe2 is the presence of Van Hove singularity (VHS) close to the Fermi level.Calandra et al. (2009); Borisenko et al. (2011) Several papers reported the successful synthesis of NbSe2 intercalated with early 3d transition metalsHe et al. (2024), while very few results have been reported for Ni intercalation, which is the focus of the present paper. When intercalating Ni in NbSe2, the nearest neighbor (NN) magnetic exchange tends to be antiferromagnetic between the planes in the case of NiNb3S6An et al. (2023). NiNb3S6 is altermagnetic with a critical temperature of 84 KAn et al. (2023) and shares the same structure as the altermagnet VNb3S6Ray et al. (2025). The NiNb4Se8 compound contains less Ni than NiNb3S6 (there are 0.25 Ni per Nb atom versus 0.33 Ni per Nb atom), but we still expect the system to be magnetic. We expect a modest critical temperature, assuming a linear dependence on Ni composition and that the magnetism is relatively insensitive to whether S or Se is present. The Ni atoms are positioned on top of the Nb atoms. The distribution of the Ni atoms is assumed to be uniform in the Density functional theory (DFT) calculations, while in the experiment, there are vacancies that will add disorder to the magnetic structure. Recently, a compound with the same space group CoNb4Se8 was found to be altermagnetic due to the antiferromagnetic coupling between the Ni plane and antiferromagnetic coupling in the ab planeRegmi et al. (2025); Vita et al. (2025). The only theoretical study about NiNb4Se8 focuses on the small unit cell (two magnetic atoms) within standard DFT, finding that the system is ferromagnetic with a low magnetic momentRamakrishna et al. (2024). We will see that we reproduce these results under the same conditions, but we expand the study to all possible magnetic phases. Experimentally, a small amount of Ni intercalation in Ni0.21NbSe2 produces a nonmagnetic system and an electron pocket at the pointWadge et al. (2025), differently from the hole pocket of the pristine NbSe2. The strong modifications of the Fermi surface, by the Ni atoms with respect to the pristine NbSe2, disrupt the CDWWadge et al. (2025).
@@
In this paper, we study the magnetic properties of NiNb4Se8, including collinear and non-collinear phases, discussing the tunability of the magnetic properties in Ni intercalated transition metal dichalcogenide NbSe2. We will demonstrate that the NiNb4Se8 transitions from a stripe antiferromagnetic phase to a ferromagnetic phase as a function of U, and discuss the possibility of a spin-glass phase. We will highlight the difference with the altermagnetic NiNb3Se6 and estimate the magnetic exchanges. The paper is structured as follows: Section 2 discusses magnetic properties with SOC, while Section 3 investigates magnetic properties including relativistic effects. In the fourth and final Section, we draw our conclusions and outlook.
@@ II Electronic and Magnetic properties without SOC
In this section, we report the structural properties and the study of the collinear magnetic phases as a function of the electronic correlations for NiNb4Se8.
@@ II.1 Structural relaxation and magnetic ground state
As it was confirmed from experimental studies for CoNb4Se8.Regmi et al. (2025), the position of the magnetic atoms tends to be on the top of Nb in transition metals intercalated NbSe2 with NiNb4Se8 stoichiometry. The experimental lattice constants of pristine bulk NbSe2 were 3.44445 Å and 12.54527 Å. The experimental lattice constants for the equivalent cell in NiNb4Se8 are 3.4540 Å and 12.428 Å, with a reduction of the in-plane lattice constant and an increase of the out-of-plane lattice constant.Wadge et al. (2025) The unit cell of NiNb4Se8 can be obtained from NbSe2 by doubling along the a and b-axis using a 221 supercell; this supercell contains two intercalated Ni atoms and allows us to simulate the magnetic phases in Fig. 1(a) and 1(b). By using a 421 supercell containing 4 magnetic atoms, we can simulate the magnetic phases reported in Fig. 1(c) and 1(d), which have two Ni atoms per plane. The Ni atoms are on a triangular lattice in the ab plane, while they are on top of each other regarding the out-of-plane connectivity. In all these 3d-intercalated transition metal dichalcogenides, we have both transition metals surrounded by 6 chalcogen atoms forming octahedra.
We use the unit cell to simulate the FM and AM phases, while we use the supercell for the AFM1 and AFM2 phases. We report the Brillouin zone (BZ) for the two space groups considered for this compound: P63/ (no. 194) for the unit cell and P21 (no. 4) for the supercell. We perform structural relaxation, constraining the volume to the experimental values. The outputs for the internal degrees of freedom are reported for the nonmagnetic phase (NM), the FM and the AM phase in Table 1, 2 and 3, respectively. We note that the internal degrees of freedom barely change from one magnetic phase to another.
@@
Given that the supercell contains 4 atoms, we can consider the ferromagnetic phase (FM) and 3 independent magnetic phases with zero magnetization. The real space pictures of these magnetic phases are reported in Fig. 1. The FM phase is reported in Fig. 1(a). We will demonstrate that the magnetic phase with in-plane ferromagnetic coupling and out-of-plane antiferromagnetic coupling is altermagnetic, we name it AM, and we report it in Fig. 1(b). The other two magnetic configurations are antiferromagnetic with Kramers degeneracy; they are illustrated in Fig. 1(c-d) and we define them as AFM1 and AFM2, respectively. AFM1 hosts both in-plane and out-of-plane antiferromagnetic coupling, while AFM2 hosts ferromagnetic coupling out-of-plane and antiferromagnetic coupling. Within the single layer, AFM1 and AFM2 have the same magnetic order, which is usually named as stripe phase magnetic order. We present in Fig. 2 the energy differences between these 4 magnetic configurations. In our theoretical results, even for U=0, the ground state is magnetic. Therefore, we expect the system to be experimentally magnetic at low temperatures in the case of nominal stoichiometry and ideal crystal structure. Note that the AM phase results in Fig. 2 are reported starting from U=1 eV, since at U=0 eV, the AM phase tends to converge to a nonmagnetic state. As shown in Fig. 2, the FM and AFM2 configurations are the most stable phases. The ground state depends on the value of the electronic correlations, there is a critical value UC at which we have a change of the ground state. The AFM2 phase is the ground state for UUC while the ferromagnetic phase is the ground state for UUC. In the non-relativistic case, we estimate U2.1 eV. As we will see in the next sections, the system is metallic with a large density of states at the Fermi level. The metallicity strongly screens the Coulomb repulsion, lowering its value. Since the experimental magnetic moments of metallic Ni1/3NbS2 are reproduced using GGA+U with U1 eVAn et al. (2023); Tenzin et al. (2025), we infer that AFM2 is the likely ground state of NiNb4Se8 under ideal stoichiometry and ordered Ni distribution. The out-of-plane ferromagnetic coupling between the Ni atoms in Ni0.25NbSe2 differs from the out-of-plane antiferromagnetic coupling found in Ni0.33NbS2 despite the two-dimensional material hosts being isoelectronic. The difference lies in the connectivity between the Ni atoms. The Ni atoms of different layers in Ni0.25NbSe2 are one on top of each other as reported in Fig. 3(b), while they are not in Ni0.33NbS2 as we can see in Fig. 3(a).
@@ II.2 Band structure
We calculate the band structure for the 4 magnetic phases and the nonmagnetic phase. The k-paths for the two space groups are reported in Fig. 4, while the band structures are reported in Fig. 5. We plot the results of the nonmagnetic phase at U=0 in Fig. 5(a), while we plot in Fig. 5(c) the band structure of the ferromagnetic phase for U=3 eV when it is the ground state. For the other phase, we plot the band structures at U=2 eV in Fig. 5(b,d,e).
@@ II.3 Fermi surface
We report the Fermi surfaces in the first BZ for the AFM2 and FM phases in Fig. 6. To compare the two phases, we use for both cases a 421 supercell (see Appendix). In this particular supercell, we do not find the hexagonal BZ of the pristine NbSe2, but the BZ is a parallelepiped. The AFM2 phase hosts 5 Fermi surfaces as shown in Fig. 6(a). The Fermi surfaces for spin-up and spin-down channels of the FM phase are plotted in Fig. 6(a) and 6(a), respectively. There are 4 Fermi surfaces for the spin-up and 6 for the spin-down channel. In all phases, we have hole- and electron-pockets.
Unlike NbSe2, which exhibits doubly nested hole pockets around the pointRossnagel et al. (2001)Johannes et al. (2006), significant electron pockets emerge in both the AFM2 and FM phases upon Ni intercalation, in agreement with band structure calculations. Along with hole-pockets present in all figures, the electron pockets at are visible mainly in the third panel of Fig. 6(a), in the second panel of Fig. 6(b), and in the fifth panel Fig. 6(c). The characteristic isolated triangular pockets observed in pristine NbSe2 vanish, and in certain bands, the Fermi surface evolves into a fully complex interconnected topology. Furthermore, compared to NbSe2, the Fermi surface dispersion along the direction is markedly enhanced. The latter is due to the Ni atoms, which work as bridges between the NbSe2 layers, making the system more three-dimensional.

### v2

@@ Abstract
We study the magnetic and electronic properties of Ni-intercalated NbSe2. We calculate the magnetic exchanges of NixNbSe2 ( and ) and find that the out-of-plane magnetic coupling depends on the Ni connectivity: it is ferromagnetic when Ni atoms stack on top of each other, and antiferromagnetic otherwise. Focusing on Ni0.25NbSe2, we identify a ground-state transition from a stripe antiferromagnetic phase with Kramers degeneracy to a ferromagnetic phase above a critical Coulomb interaction UC. Spin–orbit coupling lowers UC, aligns the easy axis along , and stabilizes collinear AFM and FM states over the competing 120∘ phase. Ni intercalation also strongly modifies the electronic structure, replacing the -point hole pocket of pristine NbSe2 with an electron pocket and shifting the Van Hove singularity away from the Fermi level, thereby suppressing potential instabilities. Finally, we investigate the altermagnetic phase in the broader class T0.25MX2, finding that spin–orbit effects induce orbital antiferromagnetism with weak ferromagnetism or ferrimagnetism depending on the Néel vector orientation. Our results demonstrate that Ni-intercalated NbSe2 provides a versatile platform to explore and tune multiple competing magnetic phases that lie close in energy.
@@ I Introduction
The tunability of magnetic and electronic phases in transition metal dichalcogenides (TMDs) doped or intercalated with ionic liquidsZhang et al. (2022), through self-intercalationWang et al. (2024), or with moleculesScheidt et al. (2015) has attracted significant research interest due to its potential for spintronic applications and the generation of new exotic phases. Intercalation of transition metals into TMDs modifies the electronic structure, leading to changes in the exchange interactions that govern the magnetic ordering. Focusing on the intercalation with magnetic transition metals, these materials exhibit diverse magnetic behaviors, including Weyl ferromagnetismSaika et al. (2022), chiral helimagnetismCao et al. (2020), altermagnetismVita et al. (2025); Candelora et al. (2025) and spin glass statesKong et al. (2023); Maniv et al. (2021). This depends on the choice of dopant, concentration, and external factors such as strain and temperatureHe et al. (2024). The spatial distribution and local environment of the intercalate atoms play a crucial role in defining the resultant magnetic order. The degree of hybridization between transition metal d-states and chalcogen p-states can significantly impact the strength and the nature of magnetic interactions.Liu et al. (2024) External stimuli, including electric gatingYu et al. (2015), pressureBai et al. (2022), and strain engineeringZhou et al. (2012), offer further tunability. Strain can enhance spin-orbit coupling (SOC) and modify superexchange interactionsChen et al. (2015), leading to novel magnetic phases. Similarly, electrostatic gating has been demonstrated to control magnetic order dynamically, enabling voltage-driven phase transitionsKhan (2024). These tunable properties make magnetic TMDs promising candidates for next-generation spintronic devices, including nonvolatile memories and logic applications.Nair et al. (2020); Shao et al. (2022)
@@
Several transition metal dichalcogenides host electronic instabilities such as charge density wave (CDW)Skolimowski et al. (2019); Wang et al. (2020); Cossu et al. (2020) and superconductivityXiao et al. (2022); Lei et al. (2020). In the ultrathin limit, NbSe2 hosts superconductivityUgeda et al. (2016); Xi et al. (2016) and CDWRivano et al. (2025), which can interplay with different layered stacking ordersCossu et al. (2024); Wang et al. (2022), self-intercalationYin et al. (2023) and can host topological phasesChiu et al. (2025). One of the relevant ingredients for the electronic instabilities in NbSe2 is the presence of Van Hove singularity (VHS) close to the Fermi level.Calandra et al. (2009); Borisenko et al. (2011) Several papers reported the successful synthesis of NbSe2 intercalated with early 3d transition metalsHe et al. (2024), while very few results have been reported for Ni intercalation, which is the focus of the present paper. When intercalating Ni in NbSe2, the nearest neighbor (NN) magnetic exchange tends to be antiferromagnetic between the planes in the case of Ni1/3NbS2An et al. (2023). Ni1/3NbS2 is altermagneticTenzin et al. (2025) with a critical temperature of 84 KAn et al. (2023) and shares the same structure as the proposed altermagnet V1/3NbS2Ray et al. (2025); Jungwirth et al. (2025); Ghosh et al. (2025). The Ni1/4NbSe2 compound contains less Ni than Ni1/3NbS2 (there are 0.25 Ni per Nb atom versus 0.33 Ni per Nb atom), but we still expect the system to be magnetic. We expect a modest critical temperature, assuming a linear dependence on Ni composition and that the magnetism is relatively insensitive to whether S or Se is present. The Ni atoms are positioned on top of the Nb atoms. The distribution of the Ni atoms is assumed to be uniform in density functional theory (DFT) calculations, while in experiment, there are vacancies that will add disorder to the magnetic structure, especially if we are far from the nominal stoichiometryErodici et al. (2023). Recently, a compound with the same space group ,CoNb4Se8, was found to be altermagnetic due to the antiferromagnetic coupling between the Ni plane and antiferromagnetic coupling in the ab planeRegmi et al. (2025); Vita et al. (2025). The only theoretical study about Ni0.25NbSe2 focuses on the small unit cell (two magnetic atoms) within standard DFT, finding that the system is ferromagnetic with a low magnetic momentRamakrishna et al. (2024).We reproduce these results under the same conditions, but we expand the study to all possible magnetic phases. Experimentally, a small amount of Ni intercalation in Ni0.19NbSe2 produces a low critical temperature and an electron pocket at the pointWadge et al. (2025a), differently from the hole pocket of the pristine NbSe2. The strong modifications of the Fermi surface, by the Ni atoms with respect to the pristine NbSe2, disrupt the CDWWadge et al. (2025b).
@@
In this paper, we study the magnetic properties of Ni0.25NbSe2 (also represented as NiNb4Se8 or NiNbSe2), including collinear and non-collinear phases, discussing the tunability of the magnetic properties in Ni intercalated transition metal dichalcogenide NbSe2. We will demonstrate that the Ni0.25NbSe2 transitions from a stripe antiferromagnetic phase to a ferromagnetic phase as a function of U, and discuss the possibility of a spin-glass phase. We will highlight the difference with the altermagnetic Ni1/3NbSe2 and estimate the magnetic exchanges. The paper is structured as follows: Section 2 discusses magnetic properties with SOC, while Section 3 investigates magnetic properties including relativistic effects. In the fourth and final Section, we draw our conclusions and outlook.
@@ II Electronic and Magnetic properties without SOC
In this section, we report the structural properties and the study of the collinear magnetic phases as a function of the electronic correlations for Ni0.25NbSe2.
@@ II.1 Structural relaxation and magnetic ground state
As it was confirmed from experimental studies for CoNb4Se8.Regmi et al. (2025), the position of the magnetic atoms tends to be on top of Nb in transition metals intercalated NbSe2 with Ni0.25NbSe2 stoichiometry. The experimental lattice constants of pristine bulk NbSe2 were 3.44445 Å and 12.54527 Å. The experimental lattice constants for the equivalent cell in Ni0.25NbSe2 are 3.455 Å and 12.3238 Å, with a reduction of the in-plane lattice constant and an increase of the out-of-plane lattice constant.Wadge et al. (2025b) The unit cell of Ni0.25NbSe2 can be obtained from NbSe2 by doubling along the a and b-axis using a 221 supercell; this supercell contains two intercalated Ni atoms and allows us to simulate the magnetic phases in Fig. 1(a) and 1(b). By using a 421 supercell containing 4 magnetic atoms, we can simulate the magnetic phases reported in Fig. 1(c) and 1(d), which have two Ni atoms per plane. The Ni atoms form a triangular lattice in the ab plane, while they are on top of each other regarding the out-of-plane connectivity. In all these 3d-intercalated transition metal dichalcogenides, we have both transition metals surrounded by 6 chalcogen atoms forming octahedra.
We use the unit cell to simulate the FM and AM phases, while we use a supercell for the AFM1 and AFM2 phases. We report the Brillouin zone (BZ) for the two space groups considered for this compound: P63/ (no. 194) for the unit cell and P21 (no. 4) for the supercell. We perform structural relaxation, constraining the volume to the experimental values. The outputs for the internal degrees of freedom are reported for the nonmagnetic phase (NM), the FM and the AM phase in Tables 1, 2 and 3, respectively. We note that the internal degrees of freedom barely change from one magnetic phase to another. It is worth mentioning that we have equivalent Nb and Se atoms. The Nb atoms at the 2b Wyckoff positions and the Se atoms at the 4f positions are the closest to the Ni atoms, while the Nb atoms at the 6h positions and the Se atoms at the 12k positions are the farthest.
@@
Given that the supercell contains 4 atoms, we can consider the ferromagnetic phase (FM) and 3 independent magnetic phases with zero magnetization. The real space pictures of these magnetic phases are reported in Fig. 1. The FM phase is reported in Fig. 1(a). We will demonstrate that the magnetic phase with in-plane ferromagnetic coupling and out-of-plane antiferromagnetic coupling is altermagnetic, we name it AM, and we report it in Fig. 1(b). The other two magnetic configurations are antiferromagnetic with Kramers’ degeneracy; they are illustrated in Fig. 1(c-d) and we define them as AFM1 and AFM2, respectively. AFM1 hosts both in-plane and out-of-plane antiferromagnetic coupling, while AFM2 hosts ferromagnetic coupling out-of-plane and antiferromagnetic coupling in-plane. Within the single layer, AFM1 and AFM2 have the same magnetic order, which is usually named as stripe phase magnetic order. We present in Fig. 2 the energy differences between these 4 magnetic configurations. In our theoretical results, even for U=0, the ground state is magnetic. Therefore, we expect the system to be experimentally magnetic at low temperatures in the case of nominal stoichiometry and ideal crystal structure. Note that the AM phase results in Fig. 2 are reported starting from U=1 eV, since at U=0 eV, the AM phase tends to converge to a nonmagnetic state with vanishing magnetic moments after structural relaxation of the internal degrees of freedom. As shown in Fig. 2, the FM and AFM2 configurations are the most stable phases. The ground state depends on the value of the electronic correlations; there is a critical value UC at which we have a change of the ground state. The AFM2 phase is the ground state for UUC while the ferromagnetic phase is the ground state for UUC. In the non-relativistic case, we estimate U2.1 eV. As we will see in the next sections, the system is metallic with a large density of states at the Fermi level. The metallicity strongly screens the Coulomb repulsion, lowering its value. Since the experimental magnetic moments of metallic Ni1/3NbS2 are reproduced using GGA+U with U1 eVAn et al. (2023); Tenzin et al. (2025), we infer that AFM2 is likely the ground state of Ni0.25NbSe2 under ideal stoichiometry and ordered Ni distribution.
The out-of-plane ferromagnetic coupling between the Ni atoms in Ni0.25NbSe2 differs from the out-of-plane antiferromagnetic coupling found in Ni1/3NbS2 despite the two-dimensional material hosts being isoelectronic. The difference lies in the connectivity between the Ni atoms. The Ni atoms of different layers in Ni0.25NbSe2 are one on top of each other as reported in Fig. 3(b), while they are not in Ni1/3NbS2 as we can see in Fig. 3(a). In Fig. 3(c) and Fig. 3(d), we report the schematic view of the crystal and magnetic structure of the altermagnetic phase of Ni0.25NbSe2 in the y–z and x–y planes, respectively. The green octahedra are the NbSe6 octahedra. The red and blue shadings mark the NiSe6 octahedra around the Ni sites with opposite spins, which are related by the twofold spin rotation combined with the sixfold crystal rotation and half-unit cell translation along the z-axis. These are the same symmetries involved in the NiAs crystal structure and produce non-relativistic spin-splitting in the band structureŠmejkal et al. (2022a); Cuono et al. (2023a, b); Fakhredine et al. (2023); therefore, the non-relativistic band structure will have the same symmetries as MnTe and CrSb with the spin-momentum locking of the kind B-4Šmejkal et al. (2022b). However, we will show later that the weak ferromagnetism will behave differently due to the presence of additional atoms, which lowers the symmetry of the magnetic space group.
@@ II.2 Band structure
We calculate the band structure for the 4 magnetic phases and the nonmagnetic phase. The k-paths for the two space groups are reported in Fig. 4, while the band structures are reported in Fig. 5. We plot the results of the nonmagnetic phase at U=0 in Fig. 5(a), while we plot in Fig. 5(c) the band structure of the ferromagnetic phase for U=3 eV when it is the ground state. For the other phases, we plot the band structures at U=2 eV in Fig. 5(b,d,e). We report the band structure for the value of U for which the relative phases are the ground state. There is a negligible difference in the band structure between U=2 and U=3 eV for the metallic phases where the Coulomb repulsion is screened; therefore, for small variations of U, the band structures will look similar.
@@ II.3 Fermi surface
We report the Fermi surfaces in the first BZ for the AFM2 and FM phases in Fig. 6. To compare the two phases, we use for both cases a 421 supercell (see Appendix). In this particular supercell, we do not find the hexagonal BZ of the pristine NbSe2, but the BZ is a parallelepiped. The AFM2 phase hosts 5 Fermi surfaces as shown in Fig. 6(a). In the FM phase, spin splitting lifts the degeneracy resulting in distinct spin-up and spin-down sheets that are plotted in Fig. 6(b) and 6(c), respectively. There are 4 Fermi surfaces for the spin-up and 6 for the spin-down channel. In all phases, we have hole- and electron-pockets. We can observe significant differences between the Fermi surfaces of the AFM2 and the FM phases, where it is more fragmented in the AFM2 phase, where smaller pockets appear at the boundaries. On the other hand, the Fermi surface of the FM phase shows smoother and more connected sheets.
Unlike the Fermi surface of NbSe2 presented in Fig. 7 which exhibits doubly nested hole pockets around the pointRossnagel et al. (2001)Johannes et al. (2006), significant electron pockets emerge in both the AFM2 and FM phases upon Ni intercalation in agreement with band structure calculations. Along with hole-pockets present in all figures, the electron pockets at are visible mainly in the third panel of Fig. 6(a), in the second panel of Fig. 6(b), and in the fifth panel of Fig. 6(c). The characteristic isolated triangular pockets observed in pristine NbSe2 vanish, and in certain bands, the Fermi surface evolves into a fully complex interconnected topology. Furthermore, compared to NbSe2, the Fermi surface dispersion along the direction is markedly enhanced. The latter is due to the Ni atoms, which work as bridges between the NbSe2 layers, making the system more three-dimensional. Both the FM and AFM2 phase have an electron pocket at the point with the FM phase showing a slightly larger pocket. Regarding the size of the Fermi sheets, the FM phase exhibits broader sheets than the AFM phase, which can be attributed to the enhanced metallicity that is usually observed in the ferromagnetic phases compared to antiferromagnetic ones.